# Individualized and Interpretable Sleep Forecasting via a two-stage Adaptive Spatial-Temporal Model

### v1

@@ IV-A Models
We propose a spatial-temporal model with attention mechanisms as our primary architecture for sleep quality prediction, as shown in the flowchart of Figure 4. This model was specifically designed to address the unique challenges of multi-modal physiological time series data, which exhibits both local patterns (e.g., short-term heart rate variations) and long-range temporal dependencies (e.g., sleep stage transitions across nights). Our architecture integrates multiple state-of-the-art components: a multi-scale CNN module with parallel convolutional layers using kernel sizes of {3, 5, 7} plus dilated convolutions, a channel attention mechanism for automatic feature selection, bidirectional LSTM layers for capturing temporal dependencies in both directions, and dual attention mechanisms including 8-head self-attention and temporal attention for identifying crucial time points. These components work synergistically to extract meaningful patterns from complex physiological signals.
We chose this hybrid architecture based on the complementary strengths of convolutional and recurrent networks for physiological time series analysis. Multi-scale CNNs excel at extracting local temporal patterns and are computationally efficient through weight sharing, while BiLSTMs capture long-term dependencies and handle variable-length sequences naturally. The attention mechanisms address the limitation of traditional RNNs in accessing distant temporal information, while the multi-scale approach acknowledges that sleep-related features manifest at different time scales [34].
@@ IV-B Domain Adaptation
In practical scenarios, the data distribution of the training set (source domain) may differ from that of the test set (target domain) due to inter-subject variability. One can clearly find domain shifts of data from each participant due to individual aspects like gender, age, lifestyle, etc, as shown in Figure 3. Figure 3 reveals significant domain shift in our dataset through PCA visualization of 16 subjects. Each participant’s data forms distinct, minimally-overlapping clusters in the reduced feature space, with subjects distributed across different regions—for example, Subject 3 occupies the lower-left quadrant while Subject 4 clusters in the upper-right. The convex hulls emphasize this separation, with some subjects (S1, S8, S10) showing compact clusters and others (S3, S7, S14) exhibiting dispersed patterns. This substantial inter-subject variability in physiological and behavioral patterns demonstrates that traditional assumptions fail for this data, necessitating domain adaptation techniques to develop robust sleep prediction models that generalize across diverse individual profiles.
@@
Domain adaptation seeks to mitigate this issue by aligning the source and target domains, thereby enhancing the model’s generalization ability. When distributions vary significantly across participants (due to physiology, behavior, etc.), each participant can be viewed as its own domain. Training on $\mathcal{D}_{p}$, validating on $\mathcal{D}_{q}$, and testing on $\mathcal{D}_{i}$ can be interpreted as adapting from a source domain to a target domain. This viewpoint motivates the application of domain adaptation methods, which aim to reduce distributional discrepancies between training participants and a new target participant.
@@
Recognizing that physiological patterns vary significantly between individuals, we designed our framework to ensure that models trained on one group of participants can effectively generalize to new, unseen participants. We implement a two-stage domain adaptation approach to address the critical challenge of cross-participant generalization in sleep quality prediction as shown in the pseudo code from Algorithm 1. Our adaptation strategy operates in two distinct stages: training-time domain adaptation and test-time adaptation (TTA). We provide flexible configuration options allowing these stages to be used independently or in combination through four modes: ’none’, ’train-only’, ’test-only’, or ’both’, enabling systematic evaluation of each component’s contribution.
@@
We introduce the second adaptation stage during inference when our model encounters data from previously unseen participants. We implement three complementary TTA strategies. First, our entropy-based TTA leverages the model’s prediction confidence for self-supervised adaptation. We identify high-confidence predictions (using a threshold of 0.9) and employ them as pseudo-labels to fine-tune the model on the test participant’s unlabeled data. Second, our consistency-based TTA enforces prediction stability under input perturbations. We update the model to minimize discrepancies between predictions on original inputs and their augmented versions (with controlled noise levels of 0.01 and 0.02). Third, we develop temporal consistency TTA that exploits the sequential nature of sleep data by enforcing smooth predictions across time and maintaining consistency between temporally shifted sequences.
@@
The training-time adaptation ensures our models learn robust, generalizable features across the training population, mitigating overfitting to individual participants’ idiosyncratic patterns. The test-time adaptation enables personalization when deploying the model to new participants, allowing dynamic adjustment to individual physiological characteristics without requiring additional labeled data. Our experimental results validate this design through systematic comparison across adaptation modes. The pseudo code is shown above in the info-graphic Algorithm 1.
@@ IV-D Explainable AI for sleep forecast
Explainable Artificial Intelligence (XAI) enhances trust and adoption by making model predictions transparent and interpretable for healthcare professionals and patients alike [35]. We implement SHAP analysis [36] using the model-agnostic KernelExplainer to interpret our deep learning models. To handle the three-dimensional input data (samples × time steps × features), we apply temporal aggregation, using 50 background samples and 100 test samples for computational efficiency. We calculate SHAP values through the standard Shapley value formula [36]. This approach provides both feature rankings and directional insights, enabling clinicians to understand not only which factors matter but how they influence sleep quality. The framework supports personalized explanations through individual SHAP values while maintaining computational tractability through temporal aggregation and strategic sampling.
@@ IV-E Training Procedure
Our training methods consist of two optional phases. In Phase 1, the model could be trained with domain adaptation using a combined loss function $\mathcal{L}=\mathcal{L}_{\text{main}}+\alpha\cdot\mathcal{L}_{\text{domain}}$, where $\mathcal{L}_{\text{main}}=\text{RMSE}(\hat{y},y)$ represents the sleep quality prediction loss and $\mathcal{L}_{\text{domain}}=\text{CrossEntropy}(\hat{d},d)$ enforces domain-invariant feature learning across participants, with the domain weight $\alpha$ optimized during hyperparameter tuning. The training employs Adam optimizer with learning rate $10^{-3}$, which incorporates early stopping with smoothed validation loss (patience=30, min_delta=0.0001), and applies comprehensive data preprocessing including ensemble smoothing methods and automated feature selection to 15 most relevant features from the original 23 Garmin metrics. In Phase 2: we implement Test-Time Adaptation (TTA) on the held-out test participant using consistency regularization with augmented inputs, minimizing $\mathcal{L}_{\text{TTA}}=\frac{1}{3}\sum_{i,j}\text{RMSE}(\hat{y}_{i},\hat{y}_{j})$ across original and noise-augmented predictions with learning rate $10^{-4}$ for 10 epochs. Phase 3 conducts final inference on the adapted model, and this entire procedure is repeated for each participant in the LOOCV framework, ensuring robust evaluation across individual differences while the domain adaptation components help the model generalize across different participants’ physiological patterns and behavioral variations in sleep data.
@@ IV-F Hyperparameter Optimization
All models in this work were optimized using Optuna’s Tree-structured Parzen Estimator (TPE) sampling within a Leave-One-Out cross-validation (LOOCV) framework to ensure robust generalization across participants. The search space included architectural parameters: number of convolutional layers $\in\{1,2\}$, LSTM layers $\in\{1,2,3\}$, CNN hidden dimensions $\in\{16,32,64\}$, LSTM hidden dimensions $\in\{64,128,256\}$, and regularization parameters: CNN dropout rate $\in[0.1,0.5]$, LSTM dropout rate $\in[0.1,0.5]$, batch normalization toggle $\in\{\text{True},\text{False}\}$, batch size $\in\{8,16,32\}$, and domain adaptation weight $\alpha\in[0.0,1.0]$. The optimization objective minimized the average RMSE across all LOOCV folds: $\theta^{*}=\arg\min_{\theta\in\Theta}\frac{1}{N}\sum_{i=1}^{N}\text{RMSE}(\mathcal{M}_{\theta},S_{i})$, where $\mathcal{M}_{\theta}$ represents the our model parameterized by $\theta$, and $S_{i}$ denotes the $i$-th held-out participant. Early stopping with smoothed validation loss (smoothing factor $\beta=0.1$) was applied with patience of 30 epochs and minimum improvement threshold of 0.0001 to prevent overfitting during the hyperparameter search process.
@@ V-A Evaluation with Different Window Sizes
We tested five input window sizes (3, 5, 7, 9, and 11 days) in combination with five predicting window sizes (1, 3, 5, 7, 9 days) on both baseline models and our proposed adaptive spatial and temporal model. Table II summarizes the results. From the results, it can be observed that all models, including our proposed approach, generally achieve their lowest RMSE when the prediction window size is one day. Longer prediction windows are also possible with the increase in errors. Our model achieves the best performance compared to other base models and state-of-the-art time series forecasting models.
Figure 5: True values vs predictions in validation and test set for subjects.
@@ V-B Temporal Prediction Analysis and Trend Accuracy
Figure 5 presents a comparison between ground-truth sleep scores and model predictions across both validation and test sets. To maintain clarity and avoid redundancy, we display representative participants that demonstrate the range of model performance patterns observed across various participants. The visualizations reveal that our model not only accurately predicts overall sleep quality levels but also effectively captures day-to-day variations and long-term sleep pattern changes for most participants.
To comprehensively evaluate model performance, we simultaneously display raw predictions and smoothed trend lines. Raw data is shown with faded colors, while trend lines are highlighted using darker colors and thicker lines. Validation and test sets are distinguished through different color schemes: blue-orange for validation set and blue-red for test set. By calculating rolling standard deviations, we add confidence bands to each prediction using semi-transparent fills to represent prediction uncertainty ranges. This visualization method intuitively demonstrates the stability of model predictions. We separately calculate Pearson correlation coefficients for validation and test sets to evaluate the model’s ability to capture sleep quality trends: Beyond traditional regression metrics, we introduce trend direction accuracy to evaluate the model’s ability to predict sleep quality change directions.
Through trend correlation assessment, we can verify whether the model captures long-term sleep pattern changes, which is crucial for monitoring chronic sleep disorders. Trend direction accuracy directly reflects the model’s ability to predict sleep quality improvement or deterioration, providing important guidance for timely interventions. By displaying prediction range bands, clinicians can understand prediction confidence levels, facilitating more cautious decision-making. Our visualization method uses color coding (green for correct direction predictions, blue for incorrect) to display model performance at each time point, enabling quick identification of model strengths and weaknesses.
@@
Trend Correlation Performance: The trend correlation values demonstrate robust model performance, test set correlations spanning 0.794 (participant 6) to 0.932 (participant 9). The consistently high correlation values (most exceeding 0.88) indicate that our model effectively captures the underlying sleep quality patterns across diverse individuals.
Trend Direction Accuracy: The trend direction accuracy metric shows performance, with test set accuracies ranging from 67.1% (participant 1) to 86.2% (participant 9). These high accuracy rates demonstrate the model’s ability to correctly predict whether sleep quality will improve or deteriorate, which is crucial for clinical intervention timing.
@@ V-B1 Individual participant Analysis
The participant-specific patterns reveal interesting variations in model performance. Participants (1, 4, 7, 9, 11) exhibit exceptional model fit with both high trend correlations (>>0.85) and direction accuracies (>>80%). The visual alignment between predicted and actual values is particularly strong, with the model successfully capturing both short-term fluctuations and long-term trends. While still achieving respectable performance metrics, participants (4, 7, 11) show slightly more prediction variance. Participant 7, for instance, demonstrates good test set performance (90.3% direction accuracy) despite having the lowest validation correlation (0.561), suggesting successful model adaptation during testing. Participants with highly variable sleep patterns (e.g., participant 11’s dramatic fluctuations around day 100) present greater prediction challenges, yet the model maintains reasonable accuracy by correctly capturing the general trend direction even when exact values differ.
@@ V-B2 Temporal Dynamics
The temporal analysis reveals several important characteristics:
Validation vs. Test Set Transition: The gray vertical line separating validation and test sets shows generally smooth transitions in prediction quality, indicating robust model generalization. The absence of significant performance degradation in test sets validates our training methodology.
Uncertainty Quantification: The shaded regions representing prediction ranges effectively bound the actual values in most cases, with wider uncertainty bands appearing during periods of higher sleep variability. This adaptive uncertainty quantification provides valuable confidence information for clinical decision-making.
Direction Prediction Visualization: The color-coded bars at the bottom of each plot (green for correct direction, blue for incorrect) provide immediate visual feedback on prediction reliability. Notably, clusters of correct predictions often coincide with stable sleep patterns, while incorrect predictions tend to occur during rapid transitions.
@@ V-B3 Clinical Relevance
The high trend direction accuracies across participants have clinical implications:
  • Early Intervention: With direction accuracies consistently above 75%, clinicians can confidently use model predictions to identify when a patient’s sleep quality is likely to deteriorate, enabling proactive interventions.
  • Personalized Prediction: The variation in model performance across participants (e.g., participant 9’s 0.932 test correlation vs. participant 6’s 0.794) suggests that some individuals have more predictable sleep patterns, allowing for customized monitoring frequencies.
  • Trend-Based Prediction: The strong trend correlations indicate that the model captures meaningful sleep quality evolution rather than just noise.
@@ V-B4 Model Robustness
The analysis demonstrates several aspects of model robustness:
  • Consistency Across participants: Despite individual sleep pattern variations, the model maintains high performance metrics across all participants.
  • Temporal Stability: The model performs consistently across extended time periods (up to 170 days for some participants), suggesting good long-term reliability.
  • Adaptation to Patterns: The model successfully adapts to different sleep pattern types, from highly regular (participant 6) to more variable (participant 11), demonstrating flexible feature learning.
@@ V-C2 Metric-Specific Observations
The MSE plot (left panel) shows the error values ranging from near 0 to 0.25. The quadratic nature of MSE amplifies the performance gaps, making it particularly useful for identifying participants where models struggle. Our model maintains the smallest area coverage, indicating consistent superiority in minimizing large prediction errors. The MAE plot (center panel) provides a more balanced view of average prediction accuracy, with values ranging from 0 to 0.4. The linear nature of MAE shows that while absolute differences between models are smaller than in MSE, the relative performance ranking remains consistent. This suggests that our model’s advantages are not merely due to outlier handling but represent genuine improvements in average prediction accuracy. The RMSE plot (right panel) combines the sensitivity of MSE to large errors with the interpretability of MAE’s scale. Values range from 0 to 0.5, and the patterns closely mirror those in the MSE plot but with a compressed scale. This metric confirms that our model’s superior performance is robust across different error measurement approaches.

### v2

@@ IV-A Models
We propose a spatial-temporal model with attention mechanisms as our primary architecture for sleep quality prediction, as shown in the flowchart of Figure 3. This model was specifically designed to address the unique challenges of multi-modal physiological time series data, which exhibits both local patterns (e.g., short-term heart rate variations) and long-range temporal dependencies.
Our architecture integrates multiple state-of-the-art components: a multi-scale CNN module with parallel convolutional layers and dilated convolutions, a channel attention mechanism for automatic feature selection, a bidirectional temporal encoder layer for capturing temporal dependencies in both directions, and dual attention mechanisms including self-attention and temporal attention for identifying crucial time points.
IV-A1 Multi-Scale Convolutional Module
Sleep-related patterns manifest at multiple temporal scales. To capture this inherent multi-scale structure, we employ parallel convolutional branches with kernel sizes of 3, 5, and 7, corresponding to short-term (1-day), medium-term (2-day), and longer-term (3-day) patterns in our daily-aggregated data. We additionally incorporate a dilated convolution branch (kernel size 3, dilation rate 2) that extends the effective receptive field to capture weekly periodicity without increasing parameter count. The outputs from all branches are concatenated to form a rich multi-scale representation. This design differs fundamentally from standard single-scale CNNs: rather than relying on deep stacking to gradually expand the receptive field, our parallel architecture explicitly extracts patterns at physiologically meaningful time scales from the first layer, enabling more efficient and interpretable feature learning.
IV-A2 Temporal modules with Dual Attention
While the CNN module excels at extracting local patterns, sleep quality prediction requires understanding longer-range temporal dependencies—for example, how accumulated sleep quality over the past week affects tonight’s sleep. We employ a bidirectional LSTM that processes the CNN features in both forward and backward directions, enabling the model to leverage both historical context and future patterns within the input window.
Building upon the LSTM representations, we introduce a dual attention mechanism that serves two complementary purposes. The first component, an 8-head self-attention layer, models pairwise interactions between all time steps, capturing non-local dependencies such as weekend-weekday contrasts or the relationship between exercise days and subsequent recovery nights. The second component, a temporal attention layer, computes a learned weighted combination over all time steps to produce the final representation. Unlike self-attention, which models relationships, temporal attention explicitly identifies which specific days within the input window are most informative for predicting current sleep quality—effectively learning that recent nights typically matter more than distant ones, while still allowing the model to attend to earlier anomalous events when relevant. The combination of self-attention and temporal attention provides richer temporal modeling than either mechanism alone: self-attention captures how different time points relate to each other, while temporal attention determines which time points matter most for the final prediction.
@@ IV-B Domain Adaptation
Figure 2 illustrates a significant domain shift in our dataset using PCA visualization across all subjects. Each participant’s data forms distinct, minimally overlapping clusters in the reduced feature space, with subjects distributed across different regions—for example, Subject 3 occupies the lower-left quadrant while Subject 4 clusters in the upper-right. The convex hulls emphasize this separation, with some subjects (S1, S8, S10) showing compact clusters and others (S3, S7, S14) exhibiting dispersed patterns. This substantial inter-subject variability in physiological and behavioral patterns demonstrates that traditional assumptions fail for this data, necessitating domain adaptation techniques to develop robust sleep prediction models that generalize across diverse individual profiles.
@@
Training on $\mathcal{D}_{p}$, validating on $\mathcal{D}_{q}$, and testing on $\mathcal{D}_{i}$ can be interpreted as adapting from a source domain to a target domain. This viewpoint motivates the application of domain adaptation methods, which aim to reduce distributional discrepancies between training participants and a new target participant.
@@
We designed our framework to ensure that models trained on one group of participants can effectively generalize to new, unseen participants. We implement a two-stage domain adaptation approach to address the critical challenge of cross-participant generalization in sleep quality prediction, as shown in the pseudo code from Algorithm 1. Our adaptation strategy operates in two distinct stages: training-time domain adaptation and TTA. We provide flexible configuration options allowing these stages to be used independently or in combination through four modes: ’none’, ’train-only’, ’test-only’, or ’both’, enabling systematic evaluation of each component’s contribution.
@@
We introduce the second adaptation stage during inference when our model encounters data from previously unseen participants. We implement three complementary TTA strategies. First, our entropy-based TTA leverages the model’s prediction confidence for self-supervised adaptation. We identify high-confidence predictions (using a threshold of 0.9) and employ them as pseudo-labels to fine-tune the model on the test participant’s unlabeled data. Second, our consistency-based TTA enforces prediction stability under input perturbations. We update the model to minimize discrepancies between predictions on original inputs and their augmented versions (with controlled noise levels of 0.01 and 0.02). Third, we develop a temporal consistency TTA that exploits the sequential nature of sleep data by enforcing smooth predictions across time and maintaining consistency between temporally shifted sequences.
@@
The training-time adaptation ensures our models learn robust and generalizable features across the training population, mitigating overfitting to individual participants’ patterns. The test-time adaptation enables personalization when deploying the model to new participants, allowing dynamic adjustment to individual physiological characteristics without requiring additional labeled data. Our experimental results validate this design through systematic comparison across adaptation modes. The pseudo code is shown above in the info-graphic Algorithm 1.
@@ IV-D Explainable AI for sleep forecast
We implement SHAP analysis [12] using the model-agnostic KernelExplainer to interpret our deep learning models. To handle the three-dimensional input data (samples × time steps × features), we apply temporal aggregation, using 50 background samples and 100 test samples for computational efficiency. We calculate SHAP values through the standard Shapley value formula [12]. This approach provides both feature rankings and directional insights, providing an understanding not only of which factors matter but also of how they influence sleep quality. The framework supports personalized explanations through individual values while maintaining computational tractability through temporal aggregation and strategic sampling.
@@ IV-E Training Procedure
[…]
@@ IV-F Hyperparameter Optimization
[…]
@@ V-A Evaluation with Different Window Sizes
We tested five input window sizes (3, 5, 7, 9, and 11 days) in combination with five predicting window sizes (1, 3, 5, 7, and 9 days) on both baseline models and our proposed adaptive spatial and temporal model. Table II summarizes the results. From the results, it can be observed that all models, including our proposed approach, generally achieve their lowest RMSE when the prediction window size is one day. Longer prediction windows are also possible with the increase in errors. Our model achieves the best performance compared to other base models and state-of-the-art time series forecasting models.
Figure 4: True values vs predictions in validation and test set for subjects. Sleep quality prediction results across subjects representing typical patterns showing real trends (solid blue lines) compared to validation predictions (orange dashed) and test predictions (red dashed) with uncertainty bands (shaded regions). Color-coded bars indicate trend direction accuracy (green = correct, blue = incorrect).
@@ V-B Temporal Prediction Analysis and Trend Accuracy
Figure 4 presents a comparison between ground-truth sleep scores and model predictions across both validation and test sets. To maintain clarity and avoid redundancy, we display representative participants who demonstrate the range of model performance patterns observed across various participants. The visualizations reveal that our model not only accurately predicts overall sleep quality levels but also effectively captures day-to-day variations and long-term sleep pattern changes for most participants.
To comprehensively evaluate model performance, we simultaneously display raw predictions and smoothed trend lines. Raw data is shown with faded colors, while trend lines are highlighted using darker colors and thicker lines. Validation and test sets are distinguished through different color schemes: blue-orange for the validation set and blue-red for the test set. By calculating rolling standard deviations, we add confidence bands to each prediction using semi-transparent fills to represent prediction uncertainty ranges. This visualization method intuitively demonstrates the stability of model predictions. We separately calculate Pearson correlation coefficients for validation and test sets to evaluate the model’s ability to capture sleep quality trends. Beyond traditional regression metrics, we introduce trend direction accuracy to evaluate the model’s ability to predict sleep quality change directions.
Through trend correlation assessment, we can verify whether the model captures long-term sleep quality changes, which is crucial for monitoring chronic sleep disorders. Trend direction accuracy directly reflects the model’s ability to predict sleep quality improvement or deterioration, providing important guidance for timely interventions. By displaying prediction range bands, clinicians can understand prediction confidence levels, facilitating more cautious decision-making. Our visualization method uses color coding (green for correct direction predictions, blue for incorrect) to display model performance at each time point, enabling quick identification of model strengths and weaknesses.
@@
Trend Correlation Performance: The trend correlation values demonstrate robust model performance, test set correlations spanning 0.864 (participant 4) to 0.941(participant 9). The consistently high correlation values (most exceeding 0.88) indicate that our model effectively captures the underlying sleep quality patterns across diverse individuals.
Trend Direction Accuracy: The trend direction accuracy metric shows performance, with test set accuracies ranging from 68.0% (participant 1) to 92.0% (participant 7). These high accuracy rates demonstrate the model’s ability to correctly predict whether sleep quality will improve or deteriorate, which is crucial for clinical intervention timing.
@@ V-B1 Individual Participant Analysis
The participant-specific patterns reveal interesting variations in model performance. Participants (4, 7, 9) exhibit exceptional model fit with both high trend correlations (>>0.864) and direction accuracies (>>75.6%). The visual alignment between predicted and actual values is particularly strong, with the model successfully capturing both short-term fluctuations and long-term trends. While still achieving respectable performance metrics, participants (7, 9) show slightly more prediction variance. Participant 7, for instance, demonstrates good test set performance (90.2% direction accuracy) despite having the lowest validation correlation (0.569), suggesting successful model adaptation during testing. Participants with highly variable sleep patterns (e.g., participant 11’s dramatic fluctuations around day 100) present greater prediction challenges, yet the model maintains reasonable accuracy by correctly capturing the general trend direction even when exact values differ.
@@ V-B2 Temporal Dynamics
The temporal analysis reveals several important characteristics. The gray vertical line separating validation and test sets shows generally smooth transitions in prediction quality, indicating robust model generalization. The absence of significant performance degradation in test sets validates our training methodology. The shaded regions representing prediction ranges effectively bound the actual values in most cases, with wider uncertainty bands appearing during periods of higher sleep variability. This adaptive uncertainty quantification provides valuable confidence information for clinical decision-making. The color-coded bars at the bottom of each plot (green for correct direction, blue for incorrect) provide immediate visual feedback on prediction reliability. Notably, clusters of correct predictions often coincide with stable sleep patterns, while incorrect predictions tend to occur during rapid transitions.
@@ V-B3 Clinical Relevance
The high trend direction accuracies across participants have clinical implications: With high direction accuracies, clinicians can confidently use model predictions to identify when a patient’s sleep quality is likely to deteriorate, enabling proactive interventions. The variation in model performance across participants (e.g., participant 9’s 0.941 test correlation vs. participant 6’s 0.896) suggests that some individuals have more predictable sleep patterns, allowing for customized monitoring frequencies. Additionally, the strong trend correlations indicate that the model captures meaningful sleep quality evolution rather than just noise, further supporting its clinical utility for personalized sleep health management.
@@ V-B4 Model Robustness
The analysis demonstrates several aspects of model robustness. Despite individual sleep pattern variations, the model maintains high performance metrics across all participants. The model performs consistently across extended time periods (up to 170 days for some participants), suggesting good long-term reliability. The model successfully adapts to different sleep pattern types, from highly regular (participant 6) to more variable (participant 11), demonstrating flexible feature learning.
@@ V-C2 Metric-Specific Observations
The MSE plot (top panel) shows the error values ranging from near 0 to 0.25. The quadratic nature of MSE amplifies the performance gaps, making it particularly useful for identifying participants where models struggle. Our model maintains the smallest area coverage, indicating consistent superiority in minimizing large prediction errors. The MAE plot (middle panel) provides a more balanced view of average prediction accuracy, with values ranging from 0 to 0.4. The linear nature of MAE shows that while absolute differences between models are smaller than in MSE, the relative performance ranking remains consistent. This suggests that our model’s advantages are not merely due to outlier handling but represent genuine improvements in average prediction accuracy. The RMSE plot (bottom panel) combines the sensitivity of MSE to large errors with the interpretability of MAE’s scale. Values range from 0 to 0.5, and the patterns closely mirror those in the MSE plot but with a compressed scale. This metric confirms that our model’s superior performance is robust across different error measurement approaches.